OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

point slope form is \[y-y1=m(x-x1)\] m=slope

OpenStudy (anonymous):

is -3 the y1? @KlOwNlOvE

OpenStudy (anonymous):

yes but remember subtracting by a negative will make it a positive

OpenStudy (anonymous):

so which one would you pick?

OpenStudy (anonymous):

so the answer is c?

OpenStudy (anonymous):

yes